# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

### v1

@@ 7 Conclusion
Real-world time series data often comes with textual descriptions, yet prior studies have largely overlooked this modality. We identify Chronological Textual Resonance, where text embeddings exhibit periodic patterns similar to their paired time series. To leverage this insight, we propose a plug-and-play framework that transforms text representations into auxiliary variables, seamlessly integrating them into existing time series models. Extensive experiments demonstrate the state-of-the-art performance of our approach.
@@
Roadmap. In this appendix, we provide a detailed overview of our methodology and experimental setup. Appendix A outlines the complete process of frequency analysis for both time series and paired texts. Appendix B includes details on datasets, hyperparameters, evaluation metrics, and additional implementation specifics. Due to space constraints in the main text, Appendix C presents the full experimental results, including comprehensive forecasting and imputation outcomes, hyperparameter and ablation studies, efficiency evaluations, and visualizations. The table of contents is provided below for reference.
@@ Appendix A Detailed Frequency Analysis Process of Time Series with Paired Texts
Here, we provide a detailed explanation of the frequency analysis process for both the time series and their paired texts.
@@
Then, we compute the Fourier Transform (Nussbaumer, 1982; Sneddon, 1995) of $\Delta\mathbf{X}$ to analyze its frequency components. The Fourier Transform of $\Delta\mathbf{X}$ is defined as:
@@
Appendix B Experiment Details
B.1 Dataset Statistics and Details
@@
Table 4 provides a summary of the statistics for the publicly available real-world datasets. Additionally, we visualize the numerical data for each multimodal time series dataset in Figure 9. For further details, please refer to the original work that introduced these datasets and benchmarks (Liu et al., 2024a).
@@
Table 4: Overview of the numerical data in the experimental datasets (Liu et al., 2024a). Our experiments utilize 9 datasets spanning multiple domains. “Prediction Length” refers to the number of future time points to be forecasted, with each dataset including four distinct prediction settings. refers to the number of variables (or variates) in each dataset.
@@
B.2 Hyperparameters
We use Adam optimizer (Kingma & Ba, 2015) when training the neural networks. The default choices of hyperparameters in our code are provided in Table 5. For LLM-based text encoders, we initialize them using the default configurations provided by Hugging Face [^1] https://huggingface.co/. Consistent with existing works (Wu et al., 2023; Liu et al., 2024c), we apply instance normalization to standardize the time series data within each dataset.
@@ B.2 Hyperparameters
pool_type Pooling type for embeddings “avg”
init_method Initialization method for combined weights “normal”
@@ B.3 Metrics
B.4 Implementation Details
B.4.1 Code and Reproducibility
@@ B.4.2 Hardware and Environment
Appendix C Full Experiment Results
C.1 Full Forecasting Performance Comparison Visualization
To provide a comprehensive comparison of different frameworks for modeling time series with paired texts, we visualize the forecasting performance using radar plots in Figure C.1. Each subfigure corresponds to a dataset, with each axis representing a different time series model. The axes are inverted, where values closer to the center indicate worse performance, and larger areas signify better results. The results demonstrate that TaTS consistently outperforms both baselines across all datasets while maintaining compatibility with various time series models.
@@ C.1 Full Forecasting Performance Comparison Visualization
C.2 Full Forecasting Results
Due to space limitations, we provide the full results of the time series forecasting task on paired time series and text in the appendix. We conduct extensive experiments across 9 datasets using 9 existing time series models, evaluating various prediction lengths as detailed in Table 4. The complete results are presented from Table 6 to Table 14. Overall, TaTS consistently achieves the best performance across all datasets, time series models, and prediction lengths. The averaged results across all prediction lengths are summarized in the main text (Table 2). For better readability, we also visualize the performance of different frameworks using radar plots, as detailed in Appendix C.1 and Figure 10.
@@ C.2 Full Forecasting Results
C.3 Full Imputation Results
@@
C.4 Full Hyperparameter Study Results of Learning Rate
@@
C.5 Full Hyperparameter Study Results of Text Embedding Dimension
@@
C.6 Full Ablation Study Results Using Different Text Encoders
@@
C.7 Full Efficiency Results: Computational Overhead vs. Performance Gain Trade-offs

### v2

@@ 7 Conclusion
Real-world time series data often comes with textual descriptions, yet prior studies have largely overlooked this modality. We identify Chronological Textual Resonance, where text embeddings exhibit periodic patterns similar to their paired time series. To leverage this insight, we propose a plug-and-play framework that transforms text representations into auxiliary variables, seamlessly integrating them into existing time series models. Extensive experiments across various forecasting models and real-world datasets demonstrate the state-of-the-art performance of our approach.
@@
Roadmap. In this appendix, we provide a detailed overview of our methodology and experimental setup. Appendix A outlines the complete process of frequency analysis for both time series and paired texts. Appendix B discusses of preprints that are related to this work. Appendix C illustrates the TaTS algorithms for time series forecasting and imputation. Appendix D includes details on datasets, hyperparameters, evaluation metrics, and additional implementation specifics. Due to space constraints in the main text, Appendix E presents the full experimental results, including comprehensive forecasting and imputation outcomes, hyperparameter and ablation studies, efficiency evaluations, and visualizations. The table of contents is provided below for reference.
@@ Appendix A Detailed Frequency Analysis Process of Time Series with Paired Texts
Here, we provide a detailed explanation of the frequency analysis process for both the time series and their paired texts. In Proposition A.1 and A.2, we simplify the proof by analyzing each periodic component individually as a single cosine. This is without loss of generality, as the operations involved are linear. Therefore, the conclusions naturally extend to signals composed of multiple periodic components through linear superposition.
@@
Then, we compute the Fourier Transform [40, 47] of $\Delta\mathbf{X}$ to analyze its frequency components. The Fourier Transform of $\Delta\mathbf{X}$ is defined as:
@@
Appendix B Further Discussion
B.1 Concurrent Works
We discuss concurrent preprints available online that relate to this research, providing a more comprehensive understanding of existing works and ongoing efforts to integrate texts into time series modeling.
[55] introduces a benchmark (CiK) for time-series forecasting that integrates both numerical data and textual context. While both [55] and our work emphasize the integration of textual context into time series forecasting, the textual contexts in [55] are descriptions regarding the time series itself, but in our work, we do not have such a constraint.
[22] develops a hybrid forecaster that jointly predicts both time series and textual data by projecting time series to the language space and fine-tuning the pre-trained LLM. Their assumption is that though LLMs are originally for language but not for time series, they can be fine-tuned to adapt to time series.
[64] introduces a dual-adapter model for time series and textual data, but their data format is similar to [51], where the whole time series is paired with one text.
[59] proposes a novel multimodal LLM framework (TS-MLLM) that treats time series as a modality akin to images. But they are using synthetic generation and focus on NLP tasks rather than time series forecasting.
B.2 Limitations
This work focuses on revealing and quantifying Chronological Textual Resonance (CTR) and designing the TaTS framework to leverage it. However, several limitations remain. First, we do not thoroughly investigate the data construction processes that may induce CTR, such as biases introduced during text collection or the choice of contextual information. Understanding how these factors affect CTR would provide deeper insights into the robustness and generalizability of our approach. Second, we do not analyze how text embeddings are utilized at the neural level within the TaTS framework. A more detailed study on how the model learns temporal patterns from paired texts could inform improvements in architecture design. Third, the effectiveness of TaTS is influenced by the quality and relevance of paired texts. While we study cases where texts are randomly shuffled, real-world texts could have different patterns, and the model’s performance may degrade. Future work could explore methods to assess and enhance text quality or develop more robust models to handle noisy input. Lastly, while TaTS demonstrates strong empirical performance across benchmark datasets, its generalization to other types of multimodal time series or domains with fundamentally different patterns remains untested, for example, time series with paired images or audio. We leave these investigations and improvements for future work.
B.3 Ethical Impact Statement
Our work is solely focused on the technical challenge of multimodal time series and does not involve any elements that could pose ethical risks.
Appendix C Algorithms
1
Input: Time series with concurrent texts embeddings $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\}$ in the input training dataset; prediction length $H$.
Output: TaTS model parameters $\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}$.
2
3Prepare training samples of sequence length $L$ and prediction length $H$ as:
$\{\mathbf{X}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{E}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{Y}^{(i)}=\mathbf{X}_{l_{i}+L+1:l_{i}+L+H}\}_{i=1}^{n}$
Initialize time series model $\mathcal{F}(\cdot;\theta_{\text{forecast}})$; projector $\text{MLP}(\cdot;\theta_{\text{MLP}})$ for dimensionality reduction.
4while not converged do
5       for each training sample $\{\mathbf{X}^{(i)},\mathbf{E}^{(i)},\mathbf{Y}^{(i)}\}$ do
6             Map $\mathbf{E}^{(i)}$ to $\mathbf{Z}^{(i)}$: $\mathbf{Z}^{(i)}[j]=\text{MLP}(\mathbf{E}^{(i)}[j];\theta_{\text{MLP}})$
7            Compute $\mathbf{U}=[\mathbf{X}^{(i)};{(\mathbf{Z}^{(i)})}^{\intercal}]_{\text{dim=1}}$ as shown in Eq. (6).
8            Forecast: $\widehat{\mathbf{X}}^{(i)}=\mathcal{F}(\mathbf{U};\theta_{\text{forecast}})[:N]$
9            Optimize: $\operatorname*{arg\,min}_{\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}}\mathcal{L}_{\text{forecast}}(\mathbf{X}^{(i)},\widehat{\mathbf{X}}^{(i)})$ as shown in Eq. (8)
10
11
return TaTS model parameters $\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}$
Algorithm 1 Texts as Time Series for Forecasting Task
1
Input: Time series with concurrent text embeddings $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\}$ in the input training dataset; binary mask $\mathbf{M}\in\{0,1\}^{T\times N}$.
Output: TaTS model parameters $\Theta=\{\theta_{\text{impute}},\theta_{\text{MLP}}\}$.
2
3Prepare training samples with observed entries:
$\{\mathbf{X}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{E}^{(i)}=\mathbf{E}_{l_{i}+1:l_{i}+L},\mathbf{M}^{(i)}=\mathbf{M}_{l_{i}+1:l_{i}+L}\}_{i=1}^{n}$
Initialize time series imputation model $\mathcal{G}(\cdot;\theta_{\text{impute}})$; projector $\text{MLP}(\cdot;\theta_{\text{MLP}})$ for dimensionality reduction.
4while not converged do
5       for each training sample $\{\mathbf{X}^{(i)},\mathbf{E}^{(i)},\mathbf{M}^{(i)}\}$ do
6             Map $\mathbf{E}^{(i)}$ to $\mathbf{Z}^{(i)}$: $\mathbf{Z}^{(i)}[j]=\text{MLP}(\mathbf{E}^{(i)}[j];\theta_{\text{MLP}})$
7            Construct the augmented input: $\mathbf{U}=[(\mathbf{X}^{(i)}\odot\mathbf{M}^{(i)});{(\mathbf{Z}^{(i)})}^{\intercal}]_{\text{dim=1}}$
8            Impute missing values: $\widehat{\mathbf{X}}^{\text{Imputed}(i)}=\mathcal{G}(\mathbf{U};\theta_{\text{impute}})$
9            Optimize: $\operatorname*{arg\,min}_{\Theta=\{\theta_{\text{impute}},\theta_{\text{MLP}}\}}\mathcal{L}_{\text{impute}}(\mathbf{X}^{(i)},\widehat{\mathbf{X}}^{\text{Imputed}(i)})$
10
11
return TaTS model parameters $\Theta=\{\theta_{\text{impute}},\theta_{\text{MLP}}\}$
Algorithm 2 Texts as Time Series for Imputation Task
Appendix D Experiment Details
D.1 Dataset Statistics and Details
@@
Table 7 provides a summary of the statistics for the publicly available real-world datasets. Additionally, we visualize the numerical data for each multimodal time series dataset in Figure 9. For further details, please refer to the original work that introduced these datasets and benchmarks [35].
@@
Table 7: Overview of the numerical data in the experimental datasets [35]. Our experiments utilize 9 datasets spanning multiple domains. “Prediction Length" refers to the number of future time points to be forecasted, with each dataset including four distinct prediction settings. refers to the number of variables (or variates) in each dataset.
@@
D.2 Hyperparameters
We use Adam optimizer [23] when training the neural networks. The default choices of hyperparameters in our code are provided in Table 8. For LLM-based text encoders, we initialize them using the default configurations provided by Hugging Face [^1] https://huggingface.co/. Consistent with existing works [57, 37], we apply instance normalization to standardize the time series data within each dataset.
@@ D.2 Hyperparameters
pool_type Pooling type for embeddings “avg"
init_method Initialization method for combined weights “normal"
@@ D.3 Metrics
D.4 Implementation Details
D.4.1 Code and Reproducibility
@@ D.4.2 Hardware and Environment
Appendix E Full Experiment Results
E.1 Why should we leverage CTR?
Method Economy Social Good Traffic
Uniformly Random (+)(+)( + ) 5.673 2.356 2.059 1.230 1.207 0.995
Normally Random 9.284 2.878 2.926 1.374 3.163 1.511
Exponentially Random 4.521 1.960 3.724 1.564 1.141 0.911
Using 1D Text only 1.995 1.404 1.315 0.853 0.714 0.797
Parallel text provides complementary information and expert knowledge that can significantly enhance the understanding of time series data. To demonstrate the benefits of utilizing periodicity in the text modality, we present an illustrative example in the univariate forecasting task, where the goal is to use $\mathbf{X}_{1:T}=\{\mathbf{x}_{1}\}\in\mathbb{R}^{T\times 1}$ to predict the next $H$ values, $\widehat{\mathbf{X}}_{T+1:T+H}$. We first concatenate the text embeddings to form $\mathbf{E}=[e_{1};e_{2};\ldots;e_{T}]_{\text{dim=1}}\in\mathbb{R}^{d_{\text{text}}\times T}$, then replace $\mathbf{x}_{1}$ with only the first dimension of the text embeddings to leverage very partial text periodicity, $\mathbf{x}^{\prime}_{1}=(\mathbf{E}[1,:])^{\intercal}\in\mathbb{R}^{T\times 1}$. In other words, we rely solely on a single evolving dimension of the paired text features to forecast future time series values.
As shown in Table 9, leveraging just the periodicity of a single text feature significantly outperforms random time series forecasting. The random baselines mean that the forecasts are purely random according to several different distributions. These random baselines use the training data as well, for example, normally random computes the mean and standard deviation of a normal distribution to forecast the time series. These results highlight that even partial periodicity from the text modality contributes valuable insights for improving forecasting accuracy.
E.2 Full Forecasting Performance Comparison Visualization
To provide a comprehensive comparison of different frameworks for modeling time series with paired texts, we visualize the forecasting performance using radar plots in Figure E.2. Each subfigure corresponds to a dataset, with each axis representing a different time series model. The axes are inverted, where values closer to the center indicate worse performance, and larger areas signify better results. The results demonstrate that TaTS consistently outperforms both baselines across all datasets while maintaining compatibility with various time series models.
@@ E.2 Full Forecasting Performance Comparison Visualization
E.3 Full Forecasting Results
Due to space limitations, we provide the full results of the time series forecasting task on paired time series and text in the appendix. We conduct extensive experiments across 9 datasets using 9 existing time series models, evaluating various prediction lengths as detailed in Table 7. The complete results are presented from Table 10 to Table 18. Overall, TaTS consistently achieves the best performance across all datasets, time series models, and prediction lengths. The averaged results across all prediction lengths are summarized in the main text (Table 2). For better readability, we also visualize the performance of different frameworks using radar plots, as detailed in Appendix E.2 and Figure 10.
@@ E.3 Full Forecasting Results
E.4 Full Imputation Results
@@
E.5 Full Hyperparameter Study Results of Learning Rate
@@
E.6 Full Hyperparameter Study Results of Text Embedding Dimension
@@
E.7 Full Ablation Study Results Using Different Text Encoders
@@
E.8 Full Efficiency Results: Computational Overhead vs. Performance Gain Trade-offs